Subject: APO 3.5m users committee minutes, April 2013

From: Suzanne Hawley

Submitted: Thu, 2 May 2013 14:05:10 -0700 (PDT)

Message number: 1309 (previous: 1308, next: 1310 up: Index)

 		APO 3.5-m Users Committee Phonecon, 4/08/13

Attending:  Suzanne Hawley, Mark Klaene, Scott Anderson, Michael Strauss, 
John Bally, Jon Holtzman, Mary Beth Kaiser, Bill Ketzeback, and Bruce 

Absent:  Al Harper, Remy Indebetouw


User feedback and comments from institutional representatives:

o Washington (Scott Anderson) - Scott had nothing new to report.

o Chicago (Al Harper) - no report.

o New Mexico State (Jon Holtzman) - Jon had nothing new to report.

o Princeton (Michael Strauss) - Michael had nothing new to report.

o Colorado (John Bally) - John had nothing new to report.

o Virginia (Remy Indebetouw) - no report.

o Johns Hopkins (Mary Beth Kaiser) - Mary Beth had nothing new to report.


Discussion of telescope/instruments report:

Mark mentioned the seasonal return of high forest fire risk at APO, and 
the usual fortifications and precautions are being re-instituted.  A 
fairly successful public open house was held.  Of special note on 
instrument issues, the Agile external timing pulse is not working, the 
echelle shutter driver board has been redesigned, and the shutter problems 
with ECam are being worked with the vendor.  Lastly, NIC-FPS still 
experiences software aborts when NFS > 8.

Mark Klaene's detailed report for the reporting period follows:


 	      3.5-m Telescope and Instruments Highlights
 	                2/28/13 through 4/01/13

0) Overview

We've had some cloud-producing weather systems but little rain/snow during 
this period.  Seasonal snow accumulation is at 18 inches, far below the 
72-inch annual average.  March proved very windy but the impact to 
operations was minimal.  Spring is settling in, so we expect more wind and 
dust storms to increase.  We held the 3rd bi-annual Public Open House this 
period.  About 90 members of the public attended, including 18 from the 
local school, and all were given a tour of the observatory after dark. 
Three groups were able to see various objects through the 3.5-m using the 

1) Telescope

The telescope performance has been nominal this period.  We did discover 
an error in the servo parameters that has to do with magna-sensor 
corrections in a negative-going direction.  This will be adjusted during 
April's engineering time.

2) Instruments

DIS has been operating nominally.

Agile had no problems this period.

SPIcam's new ion pump controller (Gamma vacuum) upgrade seems to be 
handling vacuum spikes better.

TripleSpec operated with no problems

Echelle shutter driver board failed.  The board was an undocumented 
hodgepodge of parts on a highly modified board so Ed Leon built a new 
driver board and has the shutter working again.  Also a spare shutter was 
purchased so that we have a spare unit on the shelf, given the age of the 
current shutter.  The troublesome ECam shutter is working and a spare is 
in hand.  We are also working with Apogee Instruments to see if we can use 
their new 5-million-cycle shutter.  We have started seeing communication 
disconnects with ECam and are working with Apogee -- we do have spare 
board set on hand.

NIC-FPS is operating in shared-risk observing mode due to occasional image 
corruption issue.  NFS less than 9 is operational.  The vacuum has gone 
soft and will be addressed in the coming week with a replacement ION pump 
and vacuum servicing.


LSST spectroscopy follow-up programs:

Suzanne mentioned that a meeting on the above subject is about to be held 
at NOAO.  Rene Walterbos and Michael Strauss will be attending, as will 
others with ARC/SDSS background and interests.  Suzanne envisions that our 
3.5-m telescope could participate in LSST spectroscopic follow-up much 
like we did earlier in the SDSS-II SNe survey, as well as followup on 
other transients.  A report on what happens at this meeting will be a 
topic at our next Users Committee telecon.


New APO imager:

We discussed various comments related to the new imager proposal that was 
circulated earlier.  The comments included:

   o Jon H. - no compelling science drivers to have the new imager mounted 
on a quick-change port;

   o John B. - some concerns that the design is not unique, owning a niche 
(e.g., narrow-band specialized, and/or simultaneous multi-broadband, 

   o Scott A. - agreed; a GRIZLI-like capability is much desired;

   o Suzanne - the present design does maintain narrow-band capability;

   o Michael S. - the cost of new filters will be considerable;

   o Bruce G. - what can we say about precision photometry with the new 

   o Suzanne - essentially the new design is to be considered a 
moderate-cost replacement for the aging SPICam, and be something that we 
can design and build in-house.

   o Mary Beth K. - so far, comments heard from JHU users point toward 
optimizing for higher image quality, using some kind of adaptive optics; a 
wide field is not a driver;

   o Suzanne - there is no money for an AO component in this project, but 
we may be able to implement some kind of simple tip-tilt optics;

   o John B. - we've done some testing of commercial tip-tilt systems, and 
it appears that they would not easily adapt to our image scale;

   o [unknown] - we also want to maintain the coma-free field of view that 
the telescope provides.

Suzanne concluded by saying that Joe Huenerhoff and Bill Ketzeback will 
look into these suggestions, and make a final, revised proposal in a month 
or two.  Please send any further suggestions to Joe and Bill (Joe 
Huehnerhoff <>, Bill Ketzeback 
<>), with copies to Suzanne and Mark.

The action item attached to this topic is closed.



[open from previous meeting]:

===> ACTION:  Find users to assist with instrument documentation.

 	Status:  Open.  Jon said we need to find out what improvements to 
the user documentation are needed for new users.  Suzanne said that we 
should leave this action item as a top priority; it would be good to get a 
designated "volunteer" from each institution.  Michael said he'd be happy 
to look at the DIS user documentation.

===> ACTION:  Send new ARC partner suggestions to Suzanne and Rene.

 	Status:  Open.  Some interest being seen, and there are good 
prospects.  We are also seeing interest in ARCSAT from these institutions. 
We should especially approach ex-students and other past 3.5-m users who 
are now at new institutions.  Michael offered to contact his colleagues in 
China, where there is interest in buying telescope time.  Suzanne said 
that we need to have new partner(s) in place by 2015, and if not, the 
money for instrumentation will be hard to come by, and other cost-saving 
practices will need to be employed.

===> ACTION:  Comments on DIS Slitviewer.

 	Status:  Closed.  Mark said that we are now close to a final design.

===> ACTION:  Comments on APO Imager.

 	Status:  Closed.  See discussion above.

===> ACTION:  Ideas for LSST spectroscopy followup (meeting mid-April).

 	Status:  Closed.  See discussion above.

[new from this meeting]:



Next meeting:  The next Users Committee phonecon will be on Monday 6 May 
at 8:30 AM Pacific Time.  The agenda and other materials will be sent to 
the committee members during the preceding week.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1309 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to